If the demand curve for bread is given as Q = 300 - 16p, (5 marks) a. What is the point elasticity of demand when price is 1.50? b. What will be your pricing decision if you want to increase total revenue? To increase or decrease price? 


1
Expert's answer
2021-08-24T10:13:28-0400

a) Point elasticity of demand is given by :-


"-\\frac{dQ}{dP}\\times \\frac{P}{Q}\u2212 \ndP\ndQ\n\u200b\n \u00d7 \nQ\nP\n\u200b"

B) ​Differentiating Q with respect to p we get


"\\frac{dQ}{dp} = -16\\\\ Given \\\\p = 1.50 \ndp\ndQ\n\u200b\n =\u221216\n\u200b"

Substituting this value of p in the demand equation we get

"Q=300\u221216(1.5)=300\u221224"

"Q=276"

Therefore, elasticity

"= - (-16 \u00d7 \\frac{1.50}{276}) = 0.09 (approx.)=\u2212(\u221216\u00d7 \n276\n1.50\n\u200b\n )=0.09(approx.)"

b) When the elasticity of demand is less than one, demand is said to be inelastic. This indicates that even if the price rises significantly, the change in amount requested will be minor. As a result, raising the price is a reasonable decision to raise total revenue. As a result, the quantity requested will not decrease significantly, and the corporation will be able to generate more cash.